# A stochastic method to estimate a zero-inflated two-part mixed model for human microbiome data

### v1

@@ Abstract
Human microbiome studies based on genetic sequencing techniques produce compositional longitudinal data of the relative abundances of microbial taxa over time, allowing to understand, through mixed-effects modeling, how microbial communities evolve in response to clinical interventions, environmental changes, or disease progression. In particular, the Zero-Inflated Beta Regression (ZIBR) models jointly and over time the presence and abundance of each microbe taxon, considering the compositional nature of the data, its skewness, and the over-abundance of zeros. However, as for other complex random effects models, maximum likelihood estimation suffers from the intractability of likelihood integrals. Available estimation methods rely on log-likelihood approximation, which is prone to potential limitations such as biased estimates or unstable convergence. In this work we develop an alternative maximum likelihood estimation approach for the ZIBR model, based on the Stochastic Approximation Expectation Maximization (SAEM) algorithm. The proposed methodology allows to model unbalanced data, which is not always possible in existing approaches. We also provide estimations of the standard errors and the log-likelihood of the fitted model. The performance of the algorithm is established through simulation, and its use is demonstrated on two microbiome studies, showing its ability to detect changes in both presence and abundance of bacterial taxa over time and in response to treatment.
@@ 1 Introduction
Given its major role in human health, it is crucial to have an accurate and reliable framework for the collection and analysis of human microbiome data. Current collection methods include high-throughput sequencing technologies such as the 16S ribosomal RNA (rRNA) sequencing approach and shotgun sequencing. These benchmark procedures are at the basis of leading studies such as the Human Microbiome Project (Turnbaugh et al.,, 2007), which characterized microbiomes in five major parts of the human body: airway, skin, oral cavity, digestive tract, and vagina. They produce sequence counts that are not comparable between samples, so traditional approaches consist in normalizing by the total sequence count, producing compositional data (Tyler et al.,, 2014). If multiple observations are recorded at different points in time, an additional component is introduced, leading to the analysis of longitudinal compositional data.
Regarding data analysis, Kodikara et al., (2022) compiled some recent models developed to study longitudinal microbiome data from sample sequencing, considering those that model count data and those that model relative abundances. One of the latter is the Zero-Inflated Beta Regression (ZIBR) (Chen and Li,, 2016), which is defined as a two-stage mixed effects model which is based on the work of Ospina and Ferrari, (2012), that allows the inclusion of clinical covariates, both to explain the presence or absence of a certain bacterial taxon and, in case of presence, the influence of these covariates on the relative abundance of the taxon. It also provides a comprehensive approach to analyse longitudinal compositional microbiome data taking into account its bounded nature, skewness and the over-abundance of zeros. Since its appearance, ZIBR has been succesfully applied in several studies (Hu et al.,, 2022; D’Agata et al.,, 2019) as it is capable of treating the features above mentioned, explaining within-subject correlations and providing methods to conduct hypothesis tests on the significance of covariates.
@@
As for other complex mixed-effects models, the Maximum Likelihood (ML) estimation method for ZIBR proposed by its authors relies on approximating the log-likelihood using Gauss-Hermite quadrature and numerical optimization of this expression. However, there is evidence that in certain scenarios the estimators obtained in this way may be biased and, in the case of generalized mixed models, even be outperformed by other techniques (Handayani et al.,, 2017). Additionally, the proposed estimation method can only be used for balanced data; that is, with the same number of observations per individual. In clinical studies this is not often the case, resulting in analysis with potentially misleading conclusions (Powney et al.,, 2014). Therefore, new strategies are required to address the challenges posed by missing data (Myers,, 2000). The ZIBR model can also be thought of a particular case of the Generalized Additive Model for Location, Scale and Shape (GAMLSS, Rigby and Stasinopoulos,, 2005) and its parameters can be estimated in this framework. Although in this case the available maximum estimation strategy does allow to handle missing data, it still relies on a penalized local log-likelihood approximation which can cause problems in performing likelihood ratio tests, particularly in complex GLMMs such as the ZIBR (see Stasinopoulos et al., (2017)).
@@
All things considered, the versatile features of the ZIBR model make it a promising choice for the analysis of compositional microbiome data, despite possible drawbacks in existing estimation strategies. Therefore, for the final purpose of precisely identifying those taxa responsive to disease onset, changes in environmental conditions or specific interventions, any possible improvement in the estimation method which is able to provide more accurate estimates will amount to significant progress in the understanding of human microbiome and its relation to human health.
Following this aim, in this work we propose a new estimation framework for the ZIBR model on longitudinal compositional data based on the Stochastic Approximation EM (SAEM) algorithm (Delyon et al.,, 1999). This algorithm provides an exact maximum likelihood estimation strategy in missing data models for which the EM algorithm (Dempster et al.,, 1977) is not directly applicable because the complexity of the likelihood function does not allow for exact calculation of its conditional expectation. This would be the case of the ZIBR model. The SAEM algorithm not only preserves the good behaviour of the EM algorithm in terms of convergence, unbiasedness and monotonicity, but has also shown interesting properties in complex mixed models (Márquez et al.,, 2023; Meza et al.,, 2012) and includes procedures for statistical inference and hypothesis testing (Samson et al.,, 2007). Furthermore, it can be combined with MCMC techniques for improved performance (Kuhn and Lavielle,, 2005) and can be extended to Restricted Maximum Likelihood (REML) estimation (Meza et al.,, 2007).
@@ 2 Models and methods
In this section we describe the ZIBR model for longitudinal compositional data (Chen and Li,, 2016), we revise the foundations of the SAEM algorithm for parameter estimation by maximum likelihood as well as log-likelihood estimation through Importance Sampling, and present its extension to the ZIBR model.
2.1 The ZIBR model for longitudinal compositional data
@@
$y_{it}\sim\begin{cases}0&\mbox{with prob. }1-p_{it},\\ Beta(u_{it}\phi,(1-u_{it})\phi)&\mbox{with prob. }p_{it}\\ \end{cases}$ (1)
with $\phi>0$ and $10<u_{it},p_{it}<$ 1. These two last components are characterized by
@@
where $a_{i}$ and $b_{i}$ are individual specific intercepts, $\alpha$ and $\beta$ are vectors of regression coefficients and $X_{it}$ and $Z_{it}$ are covariates for each individual and time point. We further consider that each one of the random intercepts follows a normal distribution, independently from each other:
$a_{i}\sim N(a,\sigma^{2}_{1}),\hskip 14.22636ptb_{i}\sim N(b,\sigma^{2}_{2}).$
From Equations 1 and 2, it can be seen that the ZIBR model explicitly includes a component that is responsible for the presence of zeros in the data. It is also clear that conveniently defined covariates $X_{it}$ and $Z_{it}$ can influence both the probability of presence or absence of a bacterial taxon (through the logistic regression that defines $p_{it}$) and the magnitude of its relative abundance (through the $u_{it}$ component in the proposed Beta distribution). Furthermore, the inclusion of a random intercept allows modeling correlations in observations from the same individual. Even though it is easy to expand the definition to consider random slopes, in practice it is enough to consider just a random intercept (Min and Agresti,, 2005).
The model parameter $\theta=(\phi,a,b,\alpha,\beta,\sigma^{2}_{1},\sigma^{2}_{2})$ can be estimated by maximum likelihood. From Equations 1 and 2, the likelihood function for data $\bm{y}=\left(y_{it},1\leq i\leq N,1\leq t\leq T_{i}\right)$ is
$L(\theta;\bm{y})=\prod_{i=1}^{N}\int_{\mathbb{R}}\int_{\mathbb{R}}\prod_{t=1}^{T_{i}}(1-p_{it})^{\mathbbm{1}_{\left\{y_{it}=0\right\}}}\left[p_{it}f(y_{it};u_{it},\phi)\right]^{\mathbbm{1}_{\left\{y_{it}>0\right\}}}g\left(a_{i},b_{i}|a,\sigma^{2}_{1},b,\sigma^{2}_{2}\right)da_{i}db_{i}$ (3)
@@ 2.2 The SAEM algorithm for mixed effects models
The Stochastic Approximation Expectation-Maximization (SAEM) algorithm (Delyon et al.,, 1999) is a powerful tool for estimating population parameters in complex mixed effect models. This algorithm is applicable for the iterative computation of ML estimates in a wide variety of incomplete data statistical problems in which the Expectation step of the EM algorithm is not explicit; in particular in mixed effects models, where the individual random effects are treated as non-observed data. Let $\bm{y}=\left(y_{it},1\leq i\leq N,1\leq t\leq T_{i}\right)$ and $\bm{\varphi}=\left(\varphi_{i},1\leq i\leq N\right)$ denote the observed and non-observed data, respectively, so the complete data of the model are $(\bm{y},\bm{\varphi})$. In this case, the SAEM algorithm consists of replacing the usual E-step of EM with a stochastic approximation procedure. Given an initial point $\theta^{(0)}$, iteration $q$ of the algorithm writes:
  • Simulation (S) step: Draw a realization $\bm{\varphi}^{(q)}$ from the conditional distribution $p\left(\cdot|\;\bm{y};\theta^{(q-1)}\right)$.
  • Stochastic Approximation (SA) step: Update $s_{q}(\theta)$, the approximation of the conditional expectation $\mathbb{E}\left[\log p\left(\bm{y},\bm{\varphi}^{(q)};\theta\right)|\bm{y},\theta^{(q-1)}\right]$:
    $s_{q}(\theta)=s_{q-1}(\theta)+\gamma_{q}\left(\log p\left(\bm{y},\bm{\varphi}^{(q)};\theta\right)-s_{q-1}(\theta)\right)$
@@
  • Maximization (M) step: Update $\theta^{(q)}$ according to $\theta^{(q)}=\operatorname*{arg\,max}_{\theta}s_{q}(\theta).$
There are some important remarks on the the working details of the SAEM algorithm. In the case of complex mixed effects models, such as ZIBR, the conditional distribution of the non-observed data $p\left(\cdot|\;\bm{y};\theta^{(q-1)}\right)$ cannot be computed in closed form and simulation from it cannot be carried out directly (Kuhn and Lavielle,, 2005). However, a MCMC approach can be used in the Simulation step of the SAEM algorithm described above, consisting in applying the Metropolis-Hastings algorithm (Metropolis et al.,, 1953) with different proposal kernels, in order to approximate $p\left(\cdot|\;\bm{y};\theta^{(q-1)}\right)$ with a Markov chain with defined transition probabilities.
Also, convergence can be improved by generating more than one Markov chain or realization at simulation and by applying a Monte Carlo scheme. That is, at the Simulation step $m$ realizations $\bm{\varphi}^{(q,l)}\sim p\left(\cdot|\;\bm{y};\theta^{(q-1)}\right)$, $1\leq l\leq m$, are drawn, and in the SA step the approximation of the conditional expectation is updated as
$s_{q}(\theta)=s_{q-1}(\theta)+\gamma_{q}\left(\frac{1}{m}\sum_{l=1}^{m}\log p\left(\bm{y},\bm{\varphi}^{(q,l)};\theta\right)-s_{q-1}(\theta)\right).$
If the complete-data model belongs to the exponential family, that is, if
$\log p(\bm{y},\bm{\varphi};\theta)=-\Psi(\theta)+\langle S(\bm{y},\bm{\varphi}),\xi(\theta)\rangle$
@@
and $s_{q}(\theta)=-\Psi(\theta)+\langle F_{q},\xi(\theta)\rangle$; that is, the actualization is made only on the sufficient statistic. This scheme can be applied even to models outside the exponential family, provided that a part of the model belongs to this family. However, we cannot speak of updating a sufficient statistic, but rather of a data summary function. Under general circumstances (Delyon et al.,, 1999; Kuhn and Lavielle,, 2005), the convergence of the parameter sequence $\theta^{(q)}$ toward a (local) maximum of the likelihood $\hat{\theta}$ is guaranteed, regardless of the starting point $\theta^{(0)}$ (Celeux et al.,, 1995).
The sequence of stepsizes $\{\gamma_{q}\}_{q\in\mathbb{N}}$ is usually set to 1 during the first iterations to avoid getting stuck in local maxima. In this way the first iterations are identical to those of the Monte Carlo EM (MCEM) algorithm (Wei and Tanner,, 1990), which is known for its slow convergence rate. To avoid this scenario, in later iterations of SAEM $\gamma_{q}$ decreases to zero to force convergence with fewer iterations. Details of application of the SAEM algorithm to complex mixed-effects models can be found in Meza et al., (2012); Márquez et al., (2023); Arribas-Gil et al., (2014) and de la Cruz et al., (2024).

### v2

@@ Abstract
Human microbiome studies based on genetic sequencing techniques produce longitudinal data of the relative abundances of microbial taxa over time, allowing to analyse, through mixed-effects modeling, how microbial communities evolve in response to clinical interventions, environmental changes, or disease progression. In particular, the Zero-Inflated Beta Regression (ZIBR) models jointly and over time the presence and abundance of each microbe taxon, considering the bounded nature of the data, its skewness, and the over-abundance of zeros. However, as for other complex random effects models, maximum likelihood estimation suffers from the intractability of likelihood integrals. Available estimation methods rely on log-likelihood approximation, prone to limitations such as biased estimates or unstable convergence. In this work we develop an alternative maximum likelihood estimation approach for the ZIBR model, via the Stochastic Approximation Expectation Maximization (SAEM) algorithm. The proposed methodology can accommodate unbalanced data, not always possible in existing approaches. We also provide estimations of standard errors and log-likelihood of the fitted model. The performance of the algorithm is established through simulation, and its use is demonstrated on two microbiome studies, showing its ability to detect changes in presence and abundance of bacterial taxa over time and in response to treatment.
@@ 1 Introduction
Accurate analysis of microbiome data requires reliable collection and processing frameworks. High-throughput sequencing technologies, such as the 16S ribosomal RNA (rRNA) and shotgun sequencing, underlie large initiatives such as the Human Microbiome Project (Turnbaugh et al.,, 2007). These methods generate sequence counts whose totals depend on instrument capacity, motivating normalization and the resulting compositional data structure (Tyler et al.,, 2014). This constant-sum constraint is key for comparing taxa distributions across samples, as in differential abundance analysis. However, when studying the marginal behavior of individual taxa over time or under interventions, it is common to model their proportions as bounded outcomes in $(0,1)$. In the microbiome literature, such longitudinal proportion data are often referred to as compositional even when analyzed taxon-wise, and we adopt that terminology throughout this paper.
Regarding data analysis, Kodikara et al., (2022) compiled some recent models for longitudinal microbiome data based on counts or relative abundances. Among the latter, the Zero-Inflated Beta Regression (ZIBR) model (Chen and Li,, 2016), built on the work of Ospina and Ferrari, (2012), provides a mixed-effects two-part framework to jointly model presence/absence allowing the inclusion of clinical covariates, both to explain the presence or absence and, conditional on presence, relative abundance. It naturally accommodates boundedness, skewness, and zero inflation. Since its introduction, ZIBR has been used successfully in several studies (D’Agata et al.,, 2019; Hu et al.,, 2022), capturing within-subject correlation and enabling inference on covariate effects.
@@
As for other complex mixed-effects models, ZIBR estimation via maximum likelihood (ML) relies on Gauss–Hermite quadrature and numerical optimization. Yet such approximations may introduce bias, and alternative methods often outperform ML in complex generalized mixed models (Handayani et al.,, 2017). Moreover, the proposed ML implementation is suited only for balanced data, whereas clinical studies frequently yield unbalanced or missing observations, potentially leading to biased conclusions (Myers,, 2000; Powney et al.,, 2014). The ZIBR model can also be thought of as a particular case of the Generalized Additive Model for Location, Scale and Shape (GAMLSS, Rigby and Stasinopoulos,, 2005; Stasinopoulos et al.,, 2024), which handles missing data but still relies on penalized local-likelihood approximations that complicate likelihood ratio testing in complex GLMMs (see Stasinopoulos et al., (2017)).
@@
All things considered, the versatile features of the ZIBR model make it a promising choice for the analysis of proportion microbiome data, despite possible drawbacks in existing estimation strategies. Therefore, for the final purpose of precisely identifying those taxa responsive to disease onset, changes in environmental conditions or specific interventions, any possible improvement in the estimation method which is able to provide more accurate estimates will amount to significant progress in the understanding of human microbiome and its relation to human health. Following this aim, in this work we propose a new estimation framework for the ZIBR model on longitudinal proportion data based on the Stochastic Approximation EM (SAEM) algorithm (Delyon et al.,, 1999), an exact ML strategy for missing-data models where classical EM (Dempster et al.,, 1977; McLachlan and Krishnan,, 2008) is not directly applicable due to intractable conditional expectations, precisely the case of ZIBR. The SAEM algorithm preserves the convergence and monotonicity properties of EM, has shown strong performance in complex mixed models (Meza et al.,, 2012; Márquez et al.,, 2023), provides inference and hypothesis-testing tools (Samson et al.,, 2007), and can be enhanced via MCMC techniques (Kuhn and Lavielle,, 2005) or extended to Restricted Maximum Likelihood (REML) estimation (Meza et al.,, 2007).
@@
Notice that, although the motivation for this proposal arises from the analysis of microbiome data, the potential applications of the ZIBR model extend far beyond this context. For example, in Ocaña-Riola et al., (2021) it is used to study the technical efficiency of hospitals in Spain based on their ownership structure, while in Tang et al., (2023) it is used to analyze data on plant family coverage in South Africa, adding spatial components. These are just a few examples of its possible applications.
@@ 2 Models and methods
In this section we describe the ZIBR model for longitudinal proportion data (Chen and Li,, 2016), we revise the foundations of the SAEM algorithm for parameter estimation by maximum likelihood as well as log-likelihood estimation through Importance Sampling, and present its extension to the ZIBR model.
2.1 The ZIBR model for longitudinal proportion data
@@
with $\phi>0$ and $0<u_{it},p_{it}<1$. These two last components are characterized by
$\log{\left(\frac{p_{it}}{1-p_{it}}\right)}=a_{i}+\bm{X}_{it}^{\top}\bm{\alpha},\hskip 28.45274pt\log{\left(\frac{u_{it}}{1-u_{it}}\right)}=b_{i}+\bm{Z}_{it}^{\top}\bm{\beta},$ (2)
where $a_{i}$ and $b_{i}$ are individual specific intercepts, $\bm{\alpha}$ and $\bm{\beta}$ are vectors of regression coefficients and $\bm{X}_{it}$ and $\bm{Z}_{it}$ are covariates for each individual and time point. We further consider that each one of the random intercepts follows a normal distribution, independently from each other: $a_{i}\sim N(a,\sigma^{2}_{1}),b_{i}\sim N(b,\sigma^{2}_{2}).$ From Equations 1 and 2, it can be seen that the ZIBR model explicitly includes a component that is responsible for the presence of zeros in the data. It is also clear that conveniently defined covariates $\bm{X}_{it}$ and $\bm{Z}_{it}$ can influence both the probability of presence or absence of a bacterial taxon (through the logistic regression that defines $p_{it}$) and the magnitude of its relative abundance (through the $u_{it}$ component in the proposed Beta distribution). Furthermore, the inclusion of a random intercept allows modeling correlations in observations from the same individual. Even though it is easy to expand the definition to consider random slopes, in practice it is enough to consider just a random intercept (Min and Agresti,, 2005).
The model parameter $\bm{\theta}=(\phi,a,b,\bm{\alpha},\bm{\beta},\sigma^{2}_{1},\sigma^{2}_{2})$ can be estimated by maximum likelihood. From Equations 1 and 2, the likelihood function for data $\bm{y}=\left(y_{it},1\leq i\leq N,1\leq t\leq T_{i}\right)$ is
$L(\bm{\theta};\bm{y})=\prod_{i=1}^{N}\int_{\mathbb{R}}\int_{\mathbb{R}}\prod_{t=1}^{T_{i}}(1-p_{it})^{\mathbbm{1}_{\left\{y_{it}=0\right\}}}\left[p_{it}f(y_{it};u_{it},\phi)\right]^{\mathbbm{1}_{\left\{y_{it}>0\right\}}}g\left(a_{i},b_{i}|a,\sigma^{2}_{1},b,\sigma^{2}_{2}\right)da_{i}db_{i}$ (3)
@@ 2.2 The SAEM algorithm for mixed effects models
The Stochastic Approximation Expectation-Maximization (SAEM) algorithm (Delyon et al.,, 1999) is a powerful tool for estimating population parameters in complex mixed effect models. This algorithm is applicable for the iterative computation of ML estimates in a wide variety of incomplete data statistical problems in which the Expectation step of the EM algorithm is not explicit; in particular in mixed effects models, where the individual random effects are treated as non-observed data. Let $\bm{y}=\left(y_{it},1\leq i\leq N,1\leq t\leq T_{i}\right)$ and $\bm{\varphi}=\left(\varphi_{i},1\leq i\leq N\right)$ denote the observed and non-observed data, respectively, so the complete data of the model are $(\bm{y},\bm{\varphi})$. In this case, the SAEM algorithm consists of replacing the usual E-step of EM with a stochastic approximation procedure. Given an initial point $\bm{\theta}^{(0)}$, iteration $q$ of the algorithm writes:
  • Simulation (S) step: Draw a realization $\bm{\varphi}^{(q)}$ from the conditional distribution $p\left(\cdot|\;\bm{y};\bm{\theta}^{(q-1)}\right)$.
  • Stochastic Approximation (SA) step: Update $s_{q}(\bm{\theta})$, the approximation of the conditional expectation $\mathbb{E}\left[\log p\left(\bm{y},\bm{\varphi}^{(q)};\bm{\theta}\right)|\bm{y},\bm{\theta}^{(q-1)}\right]$:
    $s_{q}(\bm{\theta})=s_{q-1}(\bm{\theta})+\gamma_{q}\left(\log p\left(\bm{y},\bm{\varphi}^{(q)};\bm{\theta}\right)-s_{q-1}(\bm{\theta})\right)$
@@
  • Maximization (M) step: Update $\bm{\theta}^{(q)}$ according to $\bm{\theta}^{(q)}=\operatorname*{arg\,max}_{\bm{\theta}}s_{q}(\bm{\theta}).$
There are some important remarks on the the working details of the SAEM algorithm. In the case of complex mixed effects models, such as ZIBR, the conditional distribution of the non-observed data $p\left(\cdot|\;\bm{y};\bm{\theta}^{(q-1)}\right)$ cannot be computed in closed form and simulation from it cannot be carried out directly (Kuhn and Lavielle,, 2005). However, a MCMC approach can be used in the Simulation step of the SAEM algorithm described above, consisting in applying the Metropolis-Hastings algorithm (Metropolis et al.,, 1953) with different proposal kernels, in order to approximate $p\left(\cdot|\;\bm{y};\bm{\theta}^{(q-1)}\right)$ with a Markov chain with defined transition probabilities.
Also, convergence can be improved by generating more than one Markov chain or realization at simulation and by applying a Monte Carlo scheme. That is, at the Simulation step $m$ realizations $\bm{\varphi}^{(q,l)}\sim p\left(\cdot|\;\bm{y};\bm{\theta}^{(q-1)}\right)$, $1\leq l\leq m$, are drawn, and in the SA step the approximation of the conditional expectation is updated as
$s_{q}(\bm{\theta})=s_{q-1}(\bm{\theta})+\gamma_{q}\left(\frac{1}{m}\sum_{l=1}^{m}\log p\left(\bm{y},\bm{\varphi}^{(q,l)};\bm{\theta}\right)-s_{q-1}(\bm{\theta})\right).$
Let $\langle a,b\rangle$ denotes the standard inner product between vectors $a$ and $b$. If the complete-data model belongs to the exponential family, that is, if
$\log p(\bm{y},\bm{\varphi};\bm{\theta})=-\Psi(\bm{\theta})+\langle S(\bm{y},\bm{\varphi}),\xi(\bm{\theta})\rangle$
@@
given that $s_{q}(\bm{\theta})=-\Psi(\bm{\theta})+\langle F_{q}(\bm{y},\bm{\varphi}),\xi(\bm{\theta})\rangle$; that is, the actualization is made only on the sufficient statistic. This scheme can be applied even to models outside the exponential family, provided that a part of the model belongs to this family. However, we cannot speak of updating a sufficient statistic, but rather of a data summary function. Under general circumstances (Delyon et al.,, 1999; Kuhn and Lavielle,, 2005), the convergence of the parameter sequence $\bm{\theta}^{(q)}$ toward a (local) maximum of the likelihood $\hat{\bm{\theta}}$ is guaranteed, regardless of the starting point $\bm{\theta}^{(0)}$ (Celeux et al.,, 1995).
The sequence of stepsizes $\{\gamma_{q}\}_{q\in\mathbb{N}}$ is usually set to 1 during the first iterations to avoid getting stuck in local maxima. In this way the first iterations are identical to those of the Monte Carlo EM (MCEM) algorithm (Wei and Tanner,, 1990), which is known for its slow convergence rate. To avoid this scenario, in later iterations of SAEM $\gamma_{q}$ decreases to zero to force convergence with fewer iterations. Details of application of the SAEM algorithm to complex mixed-effects models can be found in Meza et al., (2012); Arribas-Gil et al., (2014); Márquez et al., (2023) and de la Cruz et al., (2024).